# Semantic Communication-Enabled Wireless Adaptive Panoramic Video Transmission

### v1

@@ Abstract
In this paper, we propose an adaptive panoramic video semantic transmission (APVST) network built on the deep joint source-channel coding (Deep JSCC) structure for the efficient end-to-end transmission of panoramic videos. The proposed APVST network can adaptively extract semantic features of panoramic frames and achieve semantic feature encoding. To achieve high spectral efficiency and save bandwidth, we propose a transmission rate control mechanism for the APVST via the entropy model and the latitude adaptive model. Besides, we take weighted-to-spherically-uniform peak signal-to-noise ratio (WS-PSNR) and weighted-to-spherically-uniform structural similarity (WS-SSIM) as distortion evaluation metrics, and propose the weight attention module to fuse the weights with the semantic features to achieve better quality of immersive experiences. Finally, we evaluate our proposed scheme on a panoramic video dataset containing 208 panoramic videos. The simulation results show that the APVST can save up to 20% and 50% on channel bandwidth cost compared with other semantic communication-based and traditional video transmission schemes.
@@
Semantic communication, panoramic video, latitude adaptive, weight attention
Figure 1: The framework of APVST
@@ I Introduction
In the vision of 6G, immersive communication will be a key application scenario where users are invited to watch 360-degree panoramic videos and interact with the virtual world [1]. To ensure the quality of experience (QoE), it needs to provide low-latency and high-quality video transmission with less than 202020\,20ms [2, 3]. However, for a 4K video ($3840\times 1920$), the user can actually view the image at a resolution of approximately $960\times 540$ because of the limitations of field of view. Therefore, it needs to transmit high-resolution panoramic video to achieve a greater immersive experience, which leads to a dramatic increase in the amount of data transmitted.
Current panoramic video transmission schemes are designed based on the traditional wireless communication system, and these schemes employ bitrate selection, resource allocation, and other strategies to achieve low-latency and high transmission performance metrics [4]. However, traditional bit-level error-free transmission methods struggle to support the transmission of massive amounts of data, such as panoramic videos. As one of the key enabling technologies for 6G, semantic communication is taken into account to enhance the transmission efficiency [5, 6, 7]. Unlike traditional bit-oriented communication, it aims to pursue higher “semantic fidelity” rather than error-free transmission of symbols. By applying neural networks, artificial intelligence enables the extraction and recovery of semantic features of information, thus reducing the amount of data transmitted and achieving semantic-level compression of image, text, speech, and other types of information [8, 9].
@@
In order to improve the transmission efficiency of panoramic videos and bring better immersive QoE, a more efficient end-to-end communication system is needed. The contributions of this paper are summarized as follows: 1) We propose an end-to-end efficient adaptive panoramic video semantic transmission network, i.e., APVST. The proposed APVST is based on the Deep JSCC structure and attention mechanism to adaptively extract semantic features from panoramic frames and achieve variable-length encoding of semantic information. 2) To achieve higher WS-PSNR and WS-SSIM, we propose a weight attention (WA) module to learn the correlation between weights and semantic features for higher-quality panoramic video transmission. 3) Since ERP causes the information redundancy on different latitudes, we propose a latitude adaptive module which can reduce the bandwidth for panoramic video transmission by combining it with the entropy model.
II THE PROPOSED APVST SCHEME
In this section, we introduce the overall transmission framework of APVST, WA module, and latitude adaptive module. Finally, the optimization process is designed.
II-A The Framework of APVST
@@
The proposed APVST framework is shown in Fig. 1, which consists of the motion link and the primary link. Referring to [10][11], based on the rich correlation information between current and reference panoramic frames, the APVST estimates the motion information $\mathbf{m}_{t}$ and transmits it to the receiver through the motion link. Based on the motion information and reference panoramic frame, the context generation network extracts the context $\mathbf{c}_{t}$ and ${\hat{\mathbf{c}}}_{t}$, respectively. The primary link adaptively converts the source frames into potential representations in the semantic feature domain with the context. To enable wireless transmission, the semantic information is encoded at variable length by the Deep JSCC encoder, while the code rate is jointly guided by the learnable latitude adaptive module and entropy model. Due to the similarity between the primary link and the motion link, the transmission process of panoramic video on the primary link will be introduced in detail.
@@
After the motion link generates $\mathbf{c}_{t}$ and ${\hat{\mathbf{c}}}_{t}$, the primary link automatically learns the correlation between the current frame $\mathbf{x}_{t}$ and $\mathbf{c}_{t}$ to remove the information redundancy. The semantic feature map $\mathbf{y}_{t}$ is extracted on the basis of $\mathbf{c}_{t}$ and encoded into the corresponding codeword sequence $\mathbf{s}_{t}$ by Deep JSCC encoding, i.e,
$\mathbf{y}_{t}=f_{\mathrm{SE}}\left(\mathbf{x}_{t}\middle|\mathbf{c}_{t}\right)\ \mathrm{and}\ \mathbf{s}_{t}=f_{\mathrm{JE}}\left(\mathbf{y}_{t}^{\prime}\right),$ (1)
@@
where $f_{\mathrm{SE}}$ and $f_{\mathrm{JE}}$ denote the function of the semantic extractor and Deep JSCC encoder in the primary link, respectively. As the input to the encoder, the weighted spatial attention feature $\mathbf{y}_{t}^{\prime}$ is obtained by the WA module. The symbols through wireless channel is ${\hat{\mathbf{s}}}_{t}=\mathbf{s}_{t}+\mathbf{n}_{t}$, where $\mathbf{n}_{t}\sim\mathcal{N}\left(0,\sigma^{2}\mathrm{I}_{k_{t}}\right)$ is the noise of primary link and $k_{t}$ represents the dimension of $\mathbf{s}_{t}$. The recovered frames are reconstructed through Deep JSCC decoding and semantic recovery, which can be expressed as
${\hat{\mathbf{y}}}_{t}=f_{\mathrm{JD}}\left({\hat{\mathbf{s}}}_{t}\right)\ \mathrm{and}\ {\hat{\mathbf{x}}}_{t}=f_{\mathrm{SR}}\left({\hat{\mathbf{y}}}_{t}\middle|{\hat{\mathbf{c}}}_{t}\right),$ (2)
@@
where $f_{\mathrm{JD}}$ and $f_{\mathrm{SR}}$ denote the function of semantic restorer and Deep JSCC decoder of primary link, respectively.
@@
The APVST network is adept at autonomously learning the inter-frame correlations in panoramic videos, which is crucial for both semantic extraction and recovery. This network is particularly designed to consider key evaluation metrics during the transmission of panoramic frames, leveraging a weighted spatial attention map to significantly enhance the immersive experience, i.e., WS-PSNR and WS-SSIM. The encoding and decoding processes are meticulously guided by the entropy model and the latitude adaptive module. These components are instrumental in recognizing and adapting to the unique attributes of panoramic videos, thereby enabling targeted automatic encoding for specific regions. By this approach, APVST can effectively reduce information redundancy and achieve the efficient transmission of panoramic videos.
II-B WA Module

### v2

# Rate Splitting Multiple Access-Enabled Adaptive Panoramic Video Semantic Transmission

@@ Abstract
In this paper, we propose an adaptive panoramic video semantic transmission (APVST) framework enabled by rate splitting multiple access (RSMA). The APVST framework consists of a semantic transmitter and receiver, utilizing a deep joint source-channel coding structure to adaptively extract and encode semantic features from panoramic frames. To achieve higher spectral efficiency and conserve bandwidth, APVST employs an entropy model and a dimension-adaptive module to control the transmission rate. Additionally, we take weighted-to-spherically-uniform peak signal-to-noise ratio (WS-PSNR) and weighted-to-spherically-uniform structural similarity (WS-SSIM) as distortion evaluation metrics for panoramic videos and design a weighted self-attention module for APVST. This module integrates weights and feature maps to enhance the quality of the immersive experience. Considering the overlap in the field of view when users watch panoramic videos, we further utilize RSMA to split the required panoramic video semantic streams into common and private messages for transmission. We propose an RSMA-enabled semantic stream transmission scheme and formulate a joint problem of latency and immersive experience quality by optimizing the allocation ratios of power, common rate, and channel bandwidth, aiming to maximize the quality of service (QoS) scores for users. To address the above problem, we propose a deep reinforcement learning algorithm based on proximal policy optimization (PPO) with high efficiency to handle dynamically changing environments. Simulation results demonstrate that our proposed APVST framework saves up to 20% and 50% of channel bandwidth compared to other semantic and traditional video transmission schemes, respectively. Moreover, our study confirms the efficiency of RSMA in panoramic video transmission, achieving performance gains of 13% and 20% compared to NOMA and OFDMA.
@@
Panoramic videos, semantic communication, deep joint source-channel coding, rate splitting multiple access, deep reinforcement learning.
@@ I Introduction
In the vision of 6G, immersive communication will be a key application scenario where users are invited to watch 360-degree panoramic videos and interact with the virtual world [1], [2]. To ensure the quality of service (QoS), it needs to provide low-latency and high-quality video transmission with less than 202020\,20ms [3], [4]. However, for a 4K video ($3840\times 1920$), the user can actually view the image at a resolution of approximately $960\times 540$ because of the limitations of the field of view (FoV) [5]. Therefore, it needs to transmit high-resolution panoramic video to achieve a greater immersive experience, which leads to a dramatic increase in the amount of data transmitted.
Current panoramic video transmission schemes are designed based on the traditional wireless communication system, and these schemes employ bitrate selection, resource allocation, and other strategies to achieve low-latency and high transmission performance metrics [6]. In the context of multiple-input multiple-output (MIMO) and orthogonal frequency-division multiple access (OFDMA), Guo et al. constructed a MIMO-OFDMA system for adaptive panoramic video transmission [7]. This system is capable of adaptively selecting the quality of tiles, transmission power, and beamforming based on the current environment. Using non-orthogonal multiple access (NOMA), Li et al. proposed a NOMA-assisted VR content transmission scheme aimed at minimizing the cost-oriented towards QoS [8]. With the introduction of rate-splitting multiple access (RSMA), which is noted for its flexibility, it is increasingly being applied in various scenarios. Mao et al., in their study on RSMA, confirmed that it achieves higher spectral efficiency and system throughput compared to orthogonal multiple access (OMA), space division multiple access (SDMA), and NOMA [9].
The core concept of RSMA is to split user messages into common and private messages, enabling the capability to partially decode interference and to treat part of the interference as noise. This contrasts with the extreme interference management strategies used in SDMA and NOMA. RSMA can automatically switch between SDMA and NOMA by adjusting the power and content of common and private streams. When the interference level is low, RSMA can operate closer to the mode of SDMA, where it allocates most resources to private data streams, leveraging the advantages of spatial separation. It minimizes inter-user interference through directed beams and spatial multiplexing, thereby enhancing the system’s spectral efficiency and throughput. Conversely, when the interference level is high, RSMA tends to adopt a strategy similar to NOMA, utilizing different power levels to distinguish signals between users [9]. For the transmission of panoramic videos, RSMA can flexibly adjust the allocation ratios of common and private messages within and outside the user’s FoV, thereby reducing resource usage while ensuring the quality of user experience. Hieu et al. developed a joint communication and computational optimization problem based on RSMA, grouping users according to their FoV and employing multicast for VR video streams [10]. Huang et al. proposed a VR streaming system assisted by intelligent reflective surfaces and RSMA [11], utilizing deep reinforcement learning (DRL) algorithms to optimize IRS phase shifts, RSMA parameters, and beamforming vectors, thus enhancing QoS.
However, while existing systems that employ RSMA enhance spectral efficiency, they also potentially increase the data transmission volume by splitting the original panoramic video streams. Against this backdrop, semantic communication has garnered widespread attention as a key enabling technology for 6G [12], [13], [14]. It significantly reduces the data transmission requirement by extracting and compressing information at the semantic level, thus enhancing the efficiency of panoramic video transmission. Compared to traditional bit-level error-free transmission, semantic communication effectively overcomes the “cliff effect” caused by declines in signal-to-noise ratio (SNR) [15]. This approach not only ensures the accurate delivery of the core content and meaning of information but also enhances communication robustness in unstable or complex network environments. Unlike traditional bit-oriented communication, semantic communication aims to pursue higher “semantic fidelity” rather than error-free transmission of symbols. By applying neural networks, artificial intelligence enables the extraction and recovery of semantic features of information, thus reducing the amount of data transmitted and achieving semantic-level compression of image, text, speech, and other types of information [16], [17]. The application of this technology is expected to resolve challenges associated with simultaneous online interactions and high-quality video transmission for multiple users, thus opening new frontiers in immersive communication [18].
@@
To improve the transmission efficiency of panoramic videos and bring better immersive experience quality, a more efficient RSMA-enabled semantic communication framework is needed. Overall, we will explore how to utilize RSMA to split semantic video stream messages, thereby further enhancing the efficiency of panoramic video transmission for users. The contributions of this paper can be summarized as follows:
  • We propose an RSMA-enabled adaptive panoramic video semantic transmission framework. This framework extracts semantic features by leveraging correlations between consecutive video frames. For such a framework, considering the user’s FoV, RSMA is employed to efficiently split the semantic stream into common and private messages, thereby facilitating effective transmission of panoramic videos.
@@
  • We design a semantic transmitter and receiver for efficient adaptive panoramic video semantic transmission, i.e., APVST. The proposed APVST is based on the Deep JSCC structure and attention mechanism to adaptively extract semantic features from panoramic frames and achieve variable-length encoding of semantic information. To achieve higher WS-PSNR and WS-SSIM, we propose a weight attention (WA) module to learn the correlation between weights and semantic features for higher-quality panoramic video transmission. Since ERP causes the information redundancy on different latitudes, we propose a latitude adaptive module which can reduce the bandwidth for panoramic video transmission by combining it with the entropy model.
  • We propose a mapping method of original FoV information to the semantic level and have validated the feasibility of this mapping through empirical testing. Furthermore, we propose an RSMA-enabled semantic stream transmission scheme, which is formulated as a joint optimization problem of latency and immersive experience quality. The optimization goal of this problem is to maximize users’ QoS scores, which are defined as a weighted combination of latency score and immersive experience quality score. To address this problem, we formulate it as a Markov Decision Process (MDP) and employ the PPO-based algorithm to guide policy towards enhancing system efficiency and users’ QoS. This approach allows us to dynamically adjust the transmission strategy to accommodate environmental changes and varying demands.
  • We conducted extensive experimental validation of the proposed scheme. Simulation results indicate that our APVST framework can save 20% and 50% in channel bandwidth cost compared to other semantic and traditional video transmission schemes, respectively. Moreover, compared to NOMA and OFDMA, RSMA-enabled APVST scheme achieves gains of 13% and 20% in the transmission of semantic streams for panoramic videos.
Figure 1: System model of RSMA-enabled adaptive panoramic video semantic transmission.
The rest of this paper is organized as follows. The system model and problem formulation are described in Section II. Section III introduces the internal structures of APVST as well as its training procedure. Details of our PPO-based approach for RSMA-enabled semantic stream transmission are described in Section IV. Experimental Results are analyzed in Section V. Conclusions are finally drawn in Section VI.
Figure 2: The framework of RSMA-enabled APVST.
II System Model and Problem Formulation
We consider a scenario involving the downlink transmission of a panoramic video, which is facilitated by RSMA and based on semantic communication technology. In this scenario, a panoramic video is transmitted to $|\mathcal{U}|$ users via a single base station (BS) with $M$-antennas. Our proposed system model is illustrated in Fig. 1. The set of users is denoted by $\mathcal{U}=\{1,2,\ldots,u,\ldots,|\mathcal{U}|\}$. Given that the transmitted content is video and considering the variability of the environment, a time period is divided into multiple time slots, represented by $\mathcal{T}=\{1,2,\ldots,t,\ldots\}$. The panoramic video frame requested by user $u$ at time slot $t$ is denoted by $\mathbf{x}_{u,t}$, which is transformed into the semantic coding information $\mathbf{s}_{u,t}$ at the BS through semantic extraction and Deep JSCC encoding. Subsequently, the semantic coding information is split into common and private streams by RSMA, enabling rate adaptation and facilitating the efficient downlink transmission of semantic information. At the receiver, the user’s head-mounted display (HMD) demodulates the semantic information stream via successive interference cancellation (SIC) and achieves the recovery of the panoramic video frame $\mathbf{\hat{x}}_{u,t}$ through Deep JSCC decoding and semantic restoration.
II-A End-to-end Transmission Process
@@
The proposed RSMA-enabled adaptive panoramic video semantic transmission framework is shown in Fig. 2. We refer to the semantic transmitter and semantic receiver together as APVST, which consists of the motion link and the primary link. Referring to [19], [20], based on the rich correlation information between current and reference panoramic frames, the APVST estimates the motion information and transmits it to the receiver. Based on the motion information and reference panoramic frame, the context generator extracts the context to enable more efficient panoramic video transmission. The primary link adaptively converts the source frames into potential representations in the semantic feature domain with the context. To enable wireless transmission, the semantic information is encoded at variable length by the Deep JSCC encoder, while the code rate is jointly guided by the learnable latitude adaptive module and entropy model.
As shown in Fig. 2, the motion link estimates the motion between $\mathbf{m}_{u,t}$ the current and reference panoramic frames $\mathbf{x}_{u,t}$ and ${\mathbf{x}}_{u,t-1}$, which can be expressed as
$\mathbf{m}_{u,t}=f_{\text{ME}}\left(\mathbf{x}_{u,t},{\mathbf{x}}_{u,t-1}\right),$ (1)
where $f_{\text{ME}}$ denotes the function of Scale-space Motion Estimation. In order for the receiver to get the context, semantic extraction is performed on the $\mathbf{m}_{u,t}$ to get the motion semantic feature $\mathbf{y}_{u,t}^{\text{ml}}$. Immediately after that, it is encoded as codeword sequence $\mathbf{s}_{u,t}^{\text{ml}}$ and fed into the channel for transmission. This process can be expressed as
$\mathbf{y}_{u,t}^{\text{ml}}=f_{\text{SE}}^{\text{ml}}\left(\mathbf{m}_{u,t}\right)\ \text{and}\ \mathbf{s}_{u,t}^{\text{ml}}=f_{\text{JE}}^{\text{ml}}\left({\mathbf{y^{\prime}}}_{u,t}^{\text{ml}}\right),$ (2)
where $f_{\text{SE}}^{\text{ml}}$ and $f_{\text{JE}}^{\text{ml}}$ denote the function of the semantic extractor and Deep JSCC encoder in the motion link, respectively. As the input to the encoder, the weighted spatial attention feature ${\mathbf{y^{\prime}}}_{u,t}^{\text{ml}}$ is obtained by the WA module of the motion link. Corresponding to the transmitter, the received codeword sequence ${\hat{\mathbf{s}}}_{u,t}^{\text{ml}}$ is decoded into motion semantic feature ${\hat{\mathbf{y}}}_{u,t}^{\text{ml}}$ at the receiver. At the same time, the semantic restorer is used to recover motion information ${\hat{\mathbf{m}}}_{u,t}$, i.e,
${\hat{\mathbf{y}}}_{u,t}^{\text{ml}}=f_{\text{JD}}^{\text{ml}}\left({\hat{\mathbf{s}}}_{u,t}^{\text{ml}}\right)\ \text{and}\ {\hat{\mathbf{m}}}_{u,t}=f_{\text{SR}}^{\text{ml}}\left({\hat{\mathbf{y}}}_{u,t}^{\text{ml}}\right),$ (3)
where $f_{\text{JD}}^{\text{ml}}$ and $f_{\text{SR}}^{\text{ml}}$ denote the function of the Deep JSCC decoder and semantic restorer in the motion link, respectively. In order to obtain context $\mathbf{c}_{u,t}$ and $\mathbf{\hat{c}}_{u,t}$ at the transmitter and receiver, respectively, the motion information ${\mathbf{m}}_{u,t}$ and ${\hat{\mathbf{m}}}_{u,t}$ is fed into the context generator by combining it with the reference frames ${\mathbf{x}}_{u,t-1}$ and ${\hat{\mathbf{x}}}_{u,t-1}$. This process can be expressed as
$\mathbf{c}_{u,t}\!=\!f_{\text{CG}}\left({\mathbf{m}}_{u,t},{\mathbf{x}}_{u,t-1}\right)\text{and}\ \mathbf{\hat{c}}_{u,t}\!=\!f_{\text{CG}}\left({\hat{\mathbf{m}}}_{u,t},{\hat{\mathbf{x}}}_{u,t-1}\right),$ (4)
where $f_{\text{CG}}$ denotes the function of the context generator.
@@
After the motion link generates $\mathbf{c}_{u,t}$ and ${\hat{\mathbf{c}}}_{u,t}$, the primary link automatically learns the correlation between the current frame $\mathbf{x}_{u,t}$ and $\mathbf{c}_{u,t}$ to remove the information redundancy. The semantic feature map $\mathbf{y}_{u,t}$ is extracted on the basis of $\mathbf{c}_{u,t}$ and encoded into the corresponding codeword sequence $\mathbf{s}_{u,t}$ by Deep JSCC encoding, i.e,
$\mathbf{y}_{u,t}=f_{\text{SE}}\left(\mathbf{x}_{u,t}\middle|\mathbf{c}_{u,t}\right)\ \text{and}\ \mathbf{s}_{u,t}=f_{\text{JE}}\left(\mathbf{y}_{u,t}^{\prime}\right),$ (5)
@@
where $f_{\text{SE}}$ and $f_{\text{JE}}$ denote the function of the semantic extractor and Deep JSCC encoder in the primary link, respectively. As the input to the encoder, the weighted spatial attention feature $\mathbf{y}_{u,t}^{\prime}$ is obtained by the WA module.
In order to achieve efficient transmission of panoramic video semantic information, we will further improve users’ QoS based on RSMA. As illustrated in Fig. 1, there are overlapping and non-overlapping regions of the content viewed by the users due to their different FoVs. The encoded semantic information streams can be decomposed into common and private streams by RSMA. The former corresponds to overlapping regions and the latter corresponds to non-overlapping regions. However, due to the complexity of the neural network, the video frames have scattered the feature information to various regions during the process of semantic extraction and Deep JSCC encoding. This also directly leads to the fact that the user’s FoV region cannot be directly mapped to the corresponding semantic coding map, which reduces the interpretability of RSMA in semantic stream transmission. In the next, the feasibility of the proposed mapping scheme will be verified by experiments.
In particular, the process of masking the semantic encoding sequence $\mathbf{s}_{u,t}$ corresponding to the panoramic frame $\mathbf{x}_{u,t}$ using multiplication coefficients is as follows:
  • The coefficients within the FoV region are set to 1, indicating that no changes are made to the information within that area.
  • The mask coefficients $\rho$ outside the FoV region are set to a negative number.
  • The coefficient matrix is subjected to consecutive average pooling downsampling to obtain the mask matrix.
  • The mask matrix is vectorized to obtain the mask vector.
  • The mask vector is multiplied by the semantic encoding sequences $\mathbf{s}_{u,t}$ of the primary link and the semantic encoding sequences $\mathbf{s}^{\text{ml}}_{u,t}$ of the motion link, respectively.
  • The reconstructed visualization of the mask and the non-mask area of the panoramic frame is observed.
As illustrated in Fig. 3, we assume the current FoV of the user is the area within the red box. It can be observed that the non-mask area (i.e., the FoV area) retains the original information of the panoramic frame, whereas the mask area loses most of the panoramic frame information. This indicates that most of the information within the FoV area is preserved at the corresponding position in the semantic encoding sequence after Deep JSCC encoding. Moreover, as $\rho$ is set increasingly smaller, the image restoration quality within the FoV area deteriorates. This also demonstrates that during the semantic communication process, a small portion of information within the FoV area exists in the semantic encoding sequence corresponding to the masked area.
Based on the above experiments and analysis, the mapping from the panoramic frame level to the semantic level of FoV information can be completed by setting the weight coefficient matrix and subjecting it to downsampling and vectorization. For example, by setting the weight values within the FoV area to 1 and the weight values in other areas to $\xi$, an FoV information map $\mathbf{I}_{u,t}$ can be obtained, where $\xi$ is an adjustable smaller weight value. By subjecting $\mathbf{I}_{u,t}$ to consecutive average pooling downsampling and vectorization, the semantic-level FoV information map $\mathbf{I}_{u,t}^{s}$ can be obtained.
(a) $\rho\!=\!1,\text{PNSR}_{\text{FoV}}\!=\!36.9$
(b) $\rho\!=\!-5,\text{PNSR}_{\text{FoV}}\!=\!32.6$
(c) $\rho\!=\!-6,\text{PNSR}_{\text{FoV}}\!=\!29.8$
(d) $\rho\!=\!-7,\text{PNSR}_{\text{FoV}}\!=\!26.5$
(e) $\rho\!=\!-9,\text{PNSR}_{\text{FoV}}\!=\!21.1$
(f) $\rho\!=\!-11,\text{PNSR}_{\text{FoV}}\!=\!18.2$
(g) $\rho\!=\!-13,\text{PNSR}_{\text{FoV}}\!=\!16.5$
(h) $\rho\!=\!-15,\text{PNSR}_{\text{FoV}}\!=\!15.1$
Figure 3: The reconstructed visualization of the mask and the non-mask area of the panoramic frame. $\text{PNSR}_{\text{FoV}}$ is the PSNR of the FoV area.
Upon completing the mapping of FoV information from panoramic videos to semantic encoding sequences, the BS strategically determines which data should be incorporated into a common message and which should be categorized under private messages. Through the meticulous processing of panoramic frames’ codeword sequences $\mathbf{s}_{u,t}$ utilizing specially designed message separators and combiners, one common semantic message stream $s_{t}^{c}$ and $|\mathcal{U}|$ unique private semantic message streams $s_{u,t}^{p}$ are formed. The concepts of “common” and “private” semantic message streams reflect the overlapping and independent characteristics of the tile content viewed by users at the semantic level. Subsequently, the common semantic message stream for all users is encoded using a corresponding codebook, while the private semantic message streams for all users are encoded using the same codebook, respectively. The coding message streams $\mathbf{z}_{t}=\left[z_{t}^{c},z_{1,t}^{p},\ldots,z_{u,t}^{p},\ldots,z_{|\mathcal{U}|,t}^{p}\right]^{T}$ are created from semantic message streams, where $z_{t}^{c}$ and $z_{u,t}^{p}$ are common coding message and private coding message, respectively. The streams are allocated corresponding power at the time slot $t$ through $\mathbf{p}_{t}=\left[\text{p}_{t}^{c},\text{p}_{1,t}^{p},\ldots,\text{p}_{u,t}^{p},\ldots,\text{p}_{|\mathcal{U}|,t}^{p}\right]$. Let $P_{{\max}}$ denote the maximum transmit power of the BS, the power constraint can be expressed as
$\text{p}_{t}^{c}+\sum_{u=1}^{\left|\mathcal{U}\right|}{\text{p}_{u,t}^{p}}\leq P_{{\max}}.$ (6)
Hence, the transmit signal ultimately at the receiver can be expressed as
$\mathbf{Z}_{t}=\sqrt{\text{p}_{t}^{c}}z_{t}^{c}+\sum_{u=1}^{\left|\mathcal{U}\right|}{\sqrt{\text{p}_{u,t}^{p}}z_{t}^{p}}.$ (7)
Due to the impact of the channel, the received signal of user $u$ at time slot $t$ is expressed as
$\hat{z}_{u,t}=\sqrt{\text{p}_{t}^{c}}\mathbf{h}_{u,t}^{H}z_{t}^{c}+\mathbf{h}_{u,t}^{H}\sum_{k=1}^{\left|\mathcal{U}\right|}{\sqrt{\text{p}_{t}^{p}}z_{t}^{p}}+n_{u,t},$ (8)
where $\mathbf{h}_{u,t}$ denotes the channel matrix of user $u$ at the time slot $t$ and $n_{u,t}$ is the corresponding additive white Gaussian noise (AWGN) following $\mathcal{C}\mathcal{N}\left(0,\sigma_{u,t}^{2}\right)$.
The receiver first decodes the common semantic message stream by treating all private semantic message streams as interference. The SINR of the common semantic message stream decoded by user $u$ at time slot $t$ is
$\text{SINR}_{u,t}^{c}=\frac{\left|\mathbf{h}_{u,t}^{H}\right|^{2}\text{p}_{t}^{c}}{\sum_{k=1}^{\left|\mathcal{U}\right|}{\left|\mathbf{h}_{u,t}^{H}\right|^{2}\text{p}_{k,t}^{p}}+\sigma_{u,t}^{2}}.$ (9)
After decoding the common semantic message stream, the receiver uses SIC to subtract it. During the decoding of their private semantic message streams, other users’ private messages are considered as interference. The common semantic message stream is disregarded in this process since it has already been decoded. The SINR of private semantic message streams decoded by user $u$ at time slot $t$ can be expressed as
$\text{SINR}_{u,t}^{p}=\frac{\left|\mathbf{h}_{u,t}^{H}\right|^{2}\text{p}_{u,t}^{p}}{\sum_{k\neq u}^{\left|\mathcal{U}\right|}{\left|\mathbf{h}_{u,t}^{H}\right|^{2}\text{p}_{k,t}^{p}}+\sigma_{u,t}^{2}}.$ (10)
According to Shannon’s theorem, ideally, the transmission rate of all information cannot exceed the Shannon limit. Let $B$ denote bandwidth. The transmission rate of decoded common semantic message stream $R_{u,t}^{c}$ and private semantic message stream $R_{u,t}^{p}$ by user $u$ at time slot $t$ can be expressed as
$R_{u,t}^{c}=B\log_{2}\left(1+\text{SINR}_{u,t}^{c}\right),$ (11)
$R_{u,t}^{p}=B\log_{2}\left(1+\text{SINR}_{u,t}^{p}\right).$ (12)
In order to ensure that all users can successfully decode the semantic codeword sequence, the total common semantic message rate cannot exceed the common semantic message rate of any user, i.e,
$R_{t}^{c}=\min_{u\in\mathcal{U}}R_{u,t}^{c}.$ (13)
The restored common semantic message stream $s_{u,t}^{c}$ contains common semantic sub-messages of $|\mathcal{U}|$ users. To maximize transmission efficiency, the BS will adaptively allocate the common message rate to each user, i.e,
$\sum_{u=1}^{|\mathcal{U}|}C_{u,t}=R_{t}^{c},$ (14)
where $C_{u,t}$ is the portion of the total semantic message rate allocated to user $u$ for $s_{u,t}^{c}$. Thus, the achievable transmission rate of user $u$ at time slot $t$ is
$R_{u,t}=R_{u,t}^{p}+C_{u,t}.$ (15)
All users perform SIC to obtain the received codeword sequence is ${\hat{\mathbf{s}}}_{u,t}$ on the received signals. The recovered frame ${\hat{\mathbf{x}}}_{u,t}$ is reconstructed through Deep JSCC decoding and semantic recovery, which can be expressed as
${\hat{\mathbf{y}}}_{u,t}=f_{\text{JD}}\left({\hat{\mathbf{s}}}_{u,t}\right)\ \text{and}\ {\hat{\mathbf{x}}}_{u,t}=f_{\text{SR}}\left({\hat{\mathbf{y}}}_{u,t}\middle|{\hat{\mathbf{c}}}_{u,t}\right),$ (16)
@@
where $f_{\text{JD}}$ and $f_{\text{SR}}$ denote the function of semantic restorer and Deep JSCC decoder of primary link, respectively.
@@
The APVST network is adept at autonomously learning the inter-frame correlations in panoramic videos, which is crucial for both semantic extraction and recovery. This network is specially designed to consider key evaluation metrics during the transmission of panoramic frames, using a weighted spatial attention map to significantly enhance the quality of the immersive experience. The encoding and decoding processes are meticulously guided by the entropy model and the latitude adaptive module, enabling targeted automatic encoding for specific regions. Additionally, the model employs RSMA for the allocation of common and private messages, achieving efficient transmission of panoramic semantic message streams and further enhancing the system capacity in panoramic video transmission scenarios.
II-B Other Enable Modules for APVST
In the description of the transmission process of the system model, we mention several key modules, including the WA module and latitude adaptive module, which are crucial for the efficient transmission of panoramic videos. In the following, we will provide a detailed explanation of the working principles of these modules and their roles within the system.
II-B1 WA module